OpenStudy (anonymous):

I GIVE MEDALS Find the fifth roots of 243(cos 300° + i sin 300°).

OpenStudy (anonymous):

Thats the question I was given though

OpenStudy (anonymous):

243(cos300+isin300) \[\left( 243 \right)^{\frac{ 1 }{ 5 }}\left( \cos 300+isin300 \right)^{\frac{ 1 }{ 5 }}\] \[=\left( 3^{5} \right)^{\frac{ 1 }{ 5 }}\left( \cos \frac{ 300 }{5 }+isin \frac{ 300 }{5 }\right) \] \[=3^{5\times \frac{ 1 }{ 5 }}\left( \cos 60+i \sin 60 \right)\] \[=3\left( \frac{ 1 }{ 2 }+i \frac{\sqrt{3} }{ 2 } \right)\] Here i have used de moivre's theorem.

OpenStudy (anonymous):

The Last part is te fifth root @surjithayer ?

OpenStudy (amistre64):

since 360/5 = 72, 245(cos 300+ i sin300) 245(cos 12+ i sin12) 245(cos 84+ i sin84) 245(cos 156+ i sin156) 245(cos 228+ i sin228)

OpenStudy (anonymous):

THANK YOU OMG

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

@amistre64 but why is is 245 and not 243?

OpenStudy (amistre64):

becuase i hit the wrong key ....

OpenStudy (anonymous):

Oh okay whew lol

OpenStudy (amistre64):

:)